Find the area of a rectangle if the height is 10 and the diagonal is 26?
Answer by nerdybill(7384)

Find the area of a rectangle if the height is 10 and the diagonal is 26?
To find the area we need to know the length.
Find the width by applying Pythagorean Theorem:
Let x = length
then
x^2 + 10^2 = 26^2
x^2 + 100 = 676
x^2 = 576
x = 24
.
area = height * length = 10*24 = 240 square units